Equation x^3 + 27x^2 + p = 0 has three different real solutions that form an arithmetic progression. Find the smallest value of parameter p.


1
Expert's answer
2021-10-04T14:08:33-0400

Suppose "a, a+d, a+2d" are the roots of the cubic equation


"x^3 + 27x^2 + p = 0"

Then


"x^3 + 27x^2 + p =(x-a)(x-(a+d))(x-(a+2d))"

We have


"a+a+d+a+2d=-\dfrac{27}{1}"

"a(a+d)+a(a+2d)+(a+d)(a+2d)=\dfrac{0}{1}"

"a(a+d)(a+2d)=-\dfrac{p}{1}"

"d=-9-a"

"a(-9)+a(-a-18)+(-9)(-a-18)=0"

"-9a-a^2-18a+9a+162=0"

"a^2+18a-162=0"

"D=(18)^2-4(1)(-162)=972"

"a=\dfrac{-18\pm\sqrt{972}}{2(1)}=-9\pm9\sqrt{3}"

"a=-9-9\sqrt{3}, d=9\sqrt{3}"

"p=-(-9-9\sqrt{3})(-9)(-9+9\sqrt{3})=-1458"

"a=-9+9\sqrt{3}, d=-9\sqrt{3}"

"p=-(-9+9\sqrt{3})(-9)(-9-9\sqrt{3})=-1458"



"p=-1458"